### v1

@@ Abstract
In this paper, we investigate the prospects and challenges of sensor suites in achieving autonomous control for flying insect robots (FIRs) weighing less than a gram. FIRs, owing to their minuscule weight and size, offer unparalleled advantages in terms of material cost and scalability. However, their size introduces considerable control challenges, notably high-speed dynamics, restricted power, and limited payload capacity. While there have been notable advancements in developing lightweight sensors, often drawing inspiration from biological systems, no sub-gram aircraft has been able to attain sustained hover without relying on feedback from external sensing such as a motion capture system. The lightest vehicle capable of sustained hover—the first level of “sensor autonomy”—is the much larger 28 g Crazyflie. Previous work reported a reduction in size of that vehicle’s avionics suite to 187 mg and 21 mW. Here, we report a further reduction in mass and power to only 78.4 mg and 15 mW. We replaced the laser rangefinder with a lighter and more efficient pressure sensor, and built a smaller optic flow sensor around a global-shutter imaging chip. A Kalman Filter (KF) fuses these measurements to estimate the state variables that are needed to control hover: pitch angle, translational velocity, and altitude. Our system achieved performance comparable to that of the Crazyflie’s estimator while in flight, with root mean squared errors of 1.573 degrees, 0.186 m/s, and 0.139 m, respectively, relative to motion capture.
@@ I INTRODUCTION
Insect-scale robots have extreme size, speed, weight, and power (SSWaP) constraints. “Sensor autonomy” including stable hovering, represents first level of autonomy in the hierarchy proposed in [11], and is the foundation upon which higher level capability is built. We impose a requirement that sensing and computation to take no more than 10% of the power required to stay in the air; any more than this and flight time is severely impacted. This factor holds for important examples of autonomous drones, such as the 1.5 kg system in [12] and the 30 g system in [13]. This suggests a power budget of $\sim$20 mW for a 150 mg Robofly (Figure 1) that consumes 100 mW to fly (about 200 mW after losses from a 50% efficient boost converter are factored in).
@@
The sensing system is very crucial for robots. Large drones can have more payloads to hold sensors onboard such as inertial measurement units (IMUs) and the global positioning system (GPS). However, GPS does not work well in indoor environments [14] and small robots are not able to hold many sensors due to constraints in size, speed, weight, and power consumption (SSWaP). There are several work done on the sensor suite for insect-scale robots  [15],  [6],  [16]. The most recent sensor suite is 187 mg and consumes 21 mW [15].
In this manuscript, we introduce a new avionics system for autonomous hovering flight that is even better tailored in mass and energy consumption for an insect-scale robot (Fig. 3) than previous work [15]. To do so, we made two key refinements to lower power and mass. First, we replaced the power-hungry laser rangefinder with a pressure sensor. In recent years these have attained sufficient precision to be used for the purpose of altitude estimation and control even for insect-scale robots. Second, we replaced the optic flow sensor with a global shutter camera and compute optic flow on a microcontroller, resulting in a reduction in weight from 97 mg to 44 mg. As in [15], this new sensor suite is able to estimate the crucial variables that are needed for controlled flight: pitch angle, translational velocity, and altitude.
@@
    A significant reduction in weight of the sensor suite to 78.4 mg which is less than half of the previous lightest sensor suite [15].
@@
    We showed that we were able to make accurate estimates of the state variables pitch angle, translational velocity, and altitude using data collected wirelessly from this sensor suite while it was mounted on a flying helicopter.
@@
    Implemented an optic flow algorithm on a microcontroller.
Subsequent sections elucidate the dynamics, sensor configurations and measurement models, and corroborate our findings with data acquired wirelessly from our sensor suite, which is then analyzed on a desktop computer.
@@ II-A Dynamics
The basic dynamics of most small hovering devices are unstable in pitch angle and position. This was true for the first robotic flies [17] and EHD thruster actuated robots [18], as well as larger aircraft such as the tailless Delfly Nimble [19]. Techniques to stabilize them passively through modification of wing configuration [20] or using air dampers [21] are either unproven or result in undesirable steady-state oscillations, respectively. To stabilize these dynamics, we will use a linearized dynamical model from which we will construct an observer of these unstable states so that they can be controlled.
@@
Figure 2 depicts forces acting on a flapping-wing robot such as a gnat robot or the Robofly, but the proposed model can apply to almost any small flying aircraft, including EHD-based thrusters [22].
Near hover, the vehicle’s dynamics are well approximated by a linear state-space model [23, 24, 21] of the form $\dot{\bm{q}}=A\bm{q}+B\bm{u}+G\bm{w}$, where $\bm{w}$ is the process noise, assumed to be roughly zero-mean Gaussian white noise with known covariance. Process noise $\bm{w}$ is included because it is a necessary consideration for the design of the estimator.
To simplify the analysis, we restrict our attention to motion in the $x$-$z$ plane; the dynamics in the $y$-$z$ plane are very similar. We choose a minimal state vector $\bm{q}=[\theta,v_{x},z]^{T}$, where $\theta$ is the pitch angle, $v_{x}$ is the vehicle’s translational velocity in the world $x$-direction, and $z$ is its vertical position or altitude (Fig. 2). When the nonlinear Euler-Lagrange equation dynamics are linearized by taking the Jacobian at $\theta=0$, thrust force $f_{t}=mg$, and $z=z_{d}$, where $z_{d}$ is the desired altitude, the dynamics matrices are given by:
$\footnotesize A=\left[\begin{matrix}0&0&0\\ g&-\frac{b}{m}&0\\ 0&0&0\end{matrix}\right],\quad B=\left[\begin{matrix}1\\ 0\\ 0\end{matrix}\right]$ (1)
, where $m$ is the vehicle’s mass, and $b$ is the vehicles translational drag proportionality constant [16]. Note that because we are concerned in this paper about only estimation, we ignore inputs from the motor, and instead use the measurements of the gyroscope $\bm{u}=[\omega_{m}]$ as the only input. This is discussed further in Section V.
@@ III Sensor Suite and Measurements
Next we consider the sensor suite and state estimator. Our interest is in sensors that are either available commercially off-the-shelf, nearly so, or can be fabricated using the same tools that have been successful building insect robots such as the Robofly [25] and Robobee [26, 1]. A further constraint is that they require extremely low power for both the sensor itself and subsequent transduction and processing. This largely eliminates sensors that emit power, such as radar, sonar, depth cameras that emit structured light, and scanning laser rangefinders [27], but permits passive sensing such as vision. The global positioning system is largely denied indoors and does not provide enough resolution anyway (1–10 m), and, further, requires significant power for signal processing. A lightweight and low-power sensing modality that can provide the necessary information to hover is “optic flow,” which is defined as the velocity of the image (in rad/s) across the camera imager [28]. In tandem with other sensors, optic flow can be used to estimate the distance to obstacles as the imager translates through space [29]. Methods for estimating optic flow include calculating luminance gradients locally [30] (Lucas-Kanade), or image-wide [31], as well as iterative search for matching blocks of pixels [32]. Distances can also be estimated with a second camera and stereo vision using line matching [33], but this requires a second camera and a long “baseline” between cameras for accuracy, which would add too much weight.
Figure 3: (a) Sensor suite (PAG7920LT optical sensor, ICM42688-P IMU, BMP388 pressure sensor) next to penny.
@@
III-A The camera and optic flow
The system employs an optical sensor, the PAG7920LT (PixArt Imaging, Inc., Taiwan), to capture $160\times 120$ QQVGA images at 100 Hz over a 32 MHz SPI connection, with which the microcontroller computes optic flow via the Lucas-Kanade Method. The reason we select this camera is that, to our knowledge, it is the only global shutter camera that provides an SPI interface to transmit the image data. Other global shutter cameras require a special interface called MIPI (mobile industry processor interface), which is not available for any microcontroller. The majority of cameras employ a “rolling shutter” in which pixel rows are captured at sequential periods of time. A global shutter camera, by contrast, captures that all pixels at the same time, which ensures an accurate optic flow measurement.
@@ III-A The camera and optic flow
$A=\left[\begin{matrix}I_{x}(q_{1})&I_{y}(q_{1})\\ I_{x}(q_{2})&I_{y}(q_{2})\\ ...&...\\ I_{x}(q_{n})&I_{y}(q_{n})\end{matrix}\right]\bm{v}=\left[\begin{matrix}V_{x}\\ V_{y}\end{matrix}\right]b=\left[\begin{matrix}-I_{t}(q_{1})\\ -I_{t}(q_{2})\\ ...\\ -I_{t}(q_{n})\end{matrix}\right]$
, where $q_{1},q_{2},\dots,q_{n}$ are the brightness readings of each of the $n$ pixels in the image, $\bm{v}$ is the image flow in pixels per frame, and $I_{x}(q_{i}),I_{y}(q_{i}),I_{t}(q_{i})$ are the image’s horizontal, vertical, and temporal (time) partial derivatives respectively, at pixel $q_{i}$. Then, the optic flow is estimated (in pixels/frame) by taking the pseudoinverse $\bm{v}=(A^{T}A)^{-1}A^{T}b$.
@@
The two most expensive operations involve computing $A^{T}A$ and $A^{T}b$. A $40\times 30$ image produces an $A$ matrix with $(40-2)*(30-2)=1064$ rows and 2 columns and a $b$ matrix with 1064 rows and 1 column. We estimate multiplying two matrices of size $k\times l$ and $l\times m$ will require $kml$ multiplications. Thus, computing $A^{T}A$ and $A^{T}b$ require 4256 and 2128 multiplications, respectively. We neglect the additional operations required to compute their product (2 multiplications) and the $2\times 2$ matrix inverse (6 multiplications and one divide), giving a total operations count of 6384.
We additionally implemented a form of sparse optic flow in which flow vectors were computed for twelve $10\times 10$ patches, which are averaged to produce a final estimate. A $10\times 10$ image produces an $A$ matrix with $(10-2)*(10-2)=64$ rows and 2 columns and a $b$ matrix with 64 rows and 1 column. Thus, computing $A^{T}A$ and $A^{T}b$ require approximately 256 and 128 multiplications, respectively. With twelve matrices of size $10\times 10$, the total number of multiplication is $12(256+128)=4608$. This is a reduction of approximately $39\%$. In the future, we will evaluate these patches individually to remove outliers, removing patches with low eigenvalues (low contrast) or significantly different eigenvalues (aperture problem).
To convert the image flow in pixels/frame to angular velocity in radians/s, we rotated the TinySense by hand and minimized the error between the optic flow measurement multiplied by a constant $\alpha$ and the gyroscope reading. This constant, which depends on the pixel density and field of view of the camera, is given by $\alpha=-0.027$ rad/pixel.
@@ III-B Pressure sensor
Altitude measurements are conducted by a Bosch BMP388 pressure sensor. This sensor estimates altitude as a proportion of the difference between the surrounding air pressure and sea level. We set it for 2$\times$ pressure oversampling and an IIR filter coefficient of 3 to maximize resolution. To compensate for variations in local pressure causing inaccurate altitude measurements, the sensor was assigned a bias equal to the average of the first 25 measurements where the sensor system was grounded; this bias was subtracted from every reading.
III-C The gyroscope
@@ IV State estimation
To implement a Kalman Filter, we must first ascertain whether our system is observable at equilibrium. To show this, we linearize the observation model for the task of hovering. Note that the gyroscope provides a low-noise measurement of angular velocity $\bm{\omega}$. In cases like this, where the sensor provides more precise information than can be derived from knowing the system’s actuated inputs (e.g. the torque applied by the aircraft’s rotors or wings), it is common practice to re-formulate the Kalman Filter to use the sensor’s readings as its "input" $\bm{u}$, rather than the actual system input. Doing this has the additional advantage that for each such measurement, one less state variable is required, reducing computation requirements and complexity. In this case, it means that the angular velocity $\omega$ need not appear among the state variables. The following is our observation model for the suite of sensors, which is:
@@
where $\Omega_{m}$ is the optic flow measurements from optic flow sensor, $z_{m}$ is the measurement from the pressure sensor, and $n_{o}$ and $n_{p}$ are the noise terms for the two sensors. We assume noise terms are zero-mean Gaussian white noise, a standard assumption in sensor noise modeling to simplify the analysis. The measurement model for the optic flow in the $x$ direction, as measured by the optic flow camera is a nonlinear function of state [6]:
@@
where $\omega_{m}$ is the angular velocity measurement from gyroscope. We can see the optic flow measurement depends on $z$. We can take the Jacobian linearization at a desired altitude $z_{d}$:
@@
With this linearization, our measurement model can be cast into the traditional Kalman Filter model, $\bm{y}=C\bm{q}+D\bm{u}+\bm{n}$, where
@@
The Kalman Filter observer estimates $\hat{\bm{q}}$ by numerically integrating the system
@@
We used a gradient descent algorithm to minimize the RMSE between the TinySense state estimates and the mocap ground-truth to tune the process noise covariance matrix $Qn$ and sensor noise covariance matrix $Rn$:
$Qn=diag(0.14^{2},0.093^{2},0.08^{2}),\quad Rn=diag(1.06^{2},0.2^{2}).$
For the process noise (disturbance) matrix $G$, we assumed that the noise enters the system as white noise disturbance adding to the derivatives of the three states $\theta$, $v_{x}$, and $z$. Based on these, the Kalman gain $K$ was computed using the lqe command in python-control [34].
$G=\left[\begin{matrix}0.72&0&0\\ 0&0.10&0\\ 0&0&20.00\end{matrix}\right],\quad K=\left[\begin{matrix}0.00204&0\\ 0.48452&0\\ 0&1.81818\end{matrix}\right].$
@@
Figure 4: Sensor measurements from the onboard gyroscope, optic flow sensor, and pressure sensor of TinySense are compared with those from the Crazyflie for three different flight experiments (a), (b), and (c). In the altitude plots, altitude is measured by the Crazyflie’s laser rangefinder, while the TinySense uses a pressure sensor; mocap data is also included for comparison.

### v2

@@ Abstract
In this paper, we introduce advances in the sensor suite of an autonomous flying insect robot (FIR) weighing less than a gram. FIRs, because of their small weight and size, offer unparalleled advantages in terms of material cost and scalability. However, their size introduces considerable control challenges, notably high-speed dynamics, restricted power, and limited payload capacity. While there have been notable advancements in developing lightweight sensors, often drawing inspiration from biological systems, no sub-gram aircraft has been able to attain sustained hover without relying on feedback from external sensing such as a motion capture system. The lightest vehicle capable of sustained hovering—the first level of “sensor autonomy”—is the much larger 28 g Crazyflie. Previous work reported a reduction in size of that vehicle’s avionics suite to 187 mg and 21 mW. Here, we report a further reduction in mass and power to only 78.4 mg and 15 mW. We replaced the laser rangefinder with a lighter and more efficient pressure sensor, and built a smaller optic flow sensor around a global-shutter imaging chip. A Kalman Filter (KF) fuses these measurements to estimate the state variables that are needed to control hover: pitch angle, translational velocity, and altitude. Our system achieved performance comparable to that of the Crazyflie’s estimator while in flight, with root mean squared errors of 1.573 deg, 0.186 m/s, and 0.136 m, respectively, relative to motion capture.
@@ I INTRODUCTION
Insect-scale robots have extreme size, speed, weight, and power (SSWaP) constraints. “Sensor autonomy” including stable hovering, represents the first level of autonomy in the hierarchy proposed in [11], and is the foundation upon which higher level capability is built. We impose a requirement that sensing and computation take no more than 10% of the power required to stay in the air; any more than this and flight time is severely impacted. This factor holds for important examples of autonomous drones, such as the 1.5 kg system in [12] and the 30 g system in [13]. This suggests a power budget of $\sim$20 mW for a 150 mg Robofly (Figure 1) that consumes 100 mW to fly (about 200 mW after losses from a 50% efficient boost converter are factored in).
@@
Large drones can carry larger sensor payloads such as light-based ranging systems (LIDAR) and the global positioning system (GPS). However, GPS does not work well in indoor environments [14] and small robots are not able to hold emissive sensors due to constraints in size, speed, weight, and power consumption (SSWaP). Recent work has introduced a 187 mg avoinics suite that consumes 21 mW [15], considered gyroscope-free avionics [6], and has investigated the effect of wing-induced vibrations [16].
In this paper, we introduce a new avionics system for autonomous hovering flight that is even better tailored in mass and energy consumption for an insect-scale robot (Fig. 3) than previous work [15]. To do so, we made two key refinements to lower power and mass. First, we replaced the power-hungry laser rangefinder with a pressure sensor. In recent years, these have attained sufficient precision to be used for the purpose of altitude estimation and control, even for insect-scale robots. Second, we replaced the optic flow sensor with a global shutter camera and computed optic flow on a microcontroller, resulting in a reduction in weight from 97 mg to 44 mg. As in [15], this new sensor suite is able to estimate the crucial variables that are needed for controlled flight: pitch angle, translational velocity, and altitude.
@@
    A reduction in weight of the sensor suite to 78.4 mg which is less than half of the previous lightest sensor suite [15].
@@
    Demonstration of accurate estimates of the state variables pitch angle, translational velocity, and altitude using data collected wirelessly from this sensor suite while mounted on a drone.
@@
    Optic flow estimation on a microcontroller.
Subsequent sections elaborate on the dynamics, sensor configurations and measurement models, and validate our findings with data acquired wirelessly from our sensor suite, which is then analyzed on a desktop computer.
@@ II-A Dynamics
The basic dynamics of most small hovering devices are unstable in pitch angle and position. This was true for the first robotic flies [17] and EHD thruster actuated robots [18], as well as larger aircraft such as the tailless Delfly Nimble [19]. Techniques to stabilize them passively through modification of wing configuration [20] or using air dampers [21] are either unproven or result in undesirable steady-state oscillations, respectively. Our approach in this work is to first use linearized dynamics and observation models from which we can construct a Kalman observer of these unstable states so that they can be controlled.
@@
Fig. 2 depicts forces acting on a flapping-wing robot such as a gnat robot or the Robofly, but the proposed model can apply to almost any small flying aircraft, including EHD-based thrusters [22]. Near hover, the vehicle’s dynamics are well approximated by a linear state-space model [23, 24, 21] of the form $\dot{\bm{q}}=A\bm{q}+B\bm{u}+G\bm{w}$, where $\bm{w}$ is the process noise, assumed to be zero-mean Gaussian white noise with known covariance. Process noise $\bm{w}$ is included because it is a necessary consideration for the design of the estimator.
To simplify the analysis, we restrict our attention to motion in the $x$-$z$ plane; the dynamics in the $y$-$z$ plane are very similar. We choose a minimal state vector $\bm{q}=[\theta,v_{x},z]^{T}$, where $\theta$ is the pitch angle (rad), $v_{x}$ is the vehicle’s translational velocity in the world $x$-direction (m/s), and $z$ is its vertical position or altitude (m) as illustrated in Fig. 2. When the nonlinear Euler-Lagrange equation dynamics are linearized by taking the Jacobian at $\theta=0$, thrust force $f_{t}=mg$ (N), and $z=z_{d}$, where $z_{d}$ is the desired altitude (m), the dynamics matrices are given by:
$A=\left[\begin{matrix}0&0&0\\ g&-\frac{b}{m}&0\\ 0&0&0\end{matrix}\right],\quad B=\left[\begin{matrix}1\\ 0\\ 0\end{matrix}\right],$ (1)
where $m$ is the vehicle’s mass (kg), and $b$ is the vehicle’s translational drag proportionality constant (Ns/m) [24]. Note that because we are concerned in this paper about only estimation, we ignore inputs from the motor, and instead use the measurements of the gyroscope $\bm{u}=[\omega_{m}]$ (rad/s) as the only input. This is discussed further in Section V.
@@ III Sensor Suite and Measurements
Next we consider the sensor suite and state estimator. Our interest is in sensors that are either available commercially off-the-shelf, nearly so, or can be fabricated using the same tools that have been successful building insect robots such as the Robofly [25] and Robobee [26, 1]. A further constraint is that they require extremely low power for both the sensor itself and subsequent transduction and processing. This largely eliminates sensors that emit power, such as radar, sonar, depth cameras that emit structured light, and scanning laser rangefinders [27], but permits passive sensing such as vision. The global positioning system is largely denied indoors and does not provide enough resolution anyway (1–10 m), and, further, requires significant power for signal processing. A lightweight and low-power sensing modality that can provide the necessary information to hover is “optic flow,” which is defined as the velocity of the image (in rad/s) across the camera imager [28, 29, 30]. In tandem with other sensors, optic flow can be used to estimate the distance to obstacles as the imager translates through space [31]. Methods for estimating optic flow include calculating luminance gradients locally [32] (Lucas-Kanade), or image-wide [33], as well as iterative search for matching blocks of pixels [34]. Distances can also be estimated with a second camera and stereo vision using line matching [35], but this requires a second camera and a long “baseline” between cameras for accuracy, which adds excessive weight.
Figure 3: Sensor suite (PAG7920LT optical sensor, ICM42688-P IMU, BMP388 pressure sensor) next to US 1 cent coin.
@@
III-A Camera and optic flow
The system employs an optical sensor, the PAG7920LT (PixArt Imaging, Inc., Taiwan), to capture $160\times 120$ QQVGA images with a $120^{\circ}$ field of view at 100 Hz over a 32 MHz SPI connection, with which the microcontroller computes optic flow via the Lucas-Kanade Method. The reason we select this camera is that, to our knowledge, it is the only global shutter camera that provides an SPI interface to transmit the image data. Other global shutter cameras require a special interface called MIPI (mobile industry processor interface), which is not available for any microcontroller. The majority of cameras employ a “rolling shutter” in which pixel rows are captured at sequential periods of time. A global shutter camera, by contrast, captures that all pixels at the same time, which ensures an accurate optic flow measurement.
@@ III-A Camera and optic flow
$J=\left[\begin{matrix}I_{x}(p_{1})&I_{y}(p_{1})\\ I_{x}(p_{2})&I_{y}(p_{2})\\ ...&...\\ I_{x}(p_{n})&I_{y}(p_{n})\end{matrix}\right],\bm{v}=\left[\begin{matrix}V_{x}\\ V_{y}\end{matrix}\right],b=\left[\begin{matrix}-I_{t}(p_{1})\\ -I_{t}(p_{2})\\ ...\\ -I_{t}(p_{n})\end{matrix}\right],$
where $p_{1},p_{2},\dots,p_{n}$ are the pixels in the image, $\bm{v}$ is the image flow in pixels per frame, and $I_{x}(p_{i}),I_{y}(p_{i}),I_{t}(p_{i})$ are the image’s horizontal, vertical, and temporal (time) partial derivatives respectively, at pixel $p_{i}$. Then, the optic flow is estimated (in pixels/frame) by taking the pseudoinverse $\bm{v}=(J^{T}J)^{-1}J^{T}b$.
@@
The two most expensive operations involve computing $J^{T}J$ and $J^{T}b$. A $40\times 30$ image produces an $J$ matrix with $(40-2)\times(30-2)=1064$ rows and 2 columns and a $b$ matrix with 1064 rows and 1 column. Multiplying two matrices of size $k\times l$ and $l\times m$ requires at worst $kml$ multiplications. Thus, computing $J^{T}J$ and $J^{T}b$ require 4256 and 2128 multiplications, respectively. We neglect the additional operations required to compute their product (2 multiplications) and the $2\times 2$ matrix inverse (6 multiplications and one divide), giving a total operations count of 6384.
We additionally implemented a form of sparse optic flow in which flow vectors were computed for twelve $10\times 10$ patches, which are averaged to produce a final estimate. A $10\times 10$ image produces an $J$ matrix with $(10-2)\times(10-2)=64$ rows and 2 columns and a $b$ matrix with 64 rows and 1 column. Thus, computing $J^{T}J$ and $J^{T}b$ require approximately 256 and 128 multiplications, respectively. With twelve matrices of size $10\times 10$, the total number of multiplication is $12\times(256+128)=4608$. This is a reduction of approximately $28\%$.
To convert the optic flow from pixels/frame to angular velocity (rad/s), we found the calibration constant $\alpha=-0.027$ (rad/pixel) by rotating the TinySense by hand and performing a linear regression against its gyroscope readings.
@@ III-B Pressure sensor
Altitude measurements were conducted by a Bosch BMP388 pressure sensor. This sensor estimates altitude as a proportion of the difference between the air pressure and sea level. We set it for 2$\times$ pressure oversampling and an IIR filter coefficient of 3 to maximize resolution. To compensate for variations in local pressure causing inaccurate altitude measurements, the sensor was assigned a bias equal to the average of the first 25 measurements where the sensor system was grounded; this bias was subtracted from every reading.
III-C Gyroscope
@@ IV State estimation
To implement a Kalman Filter, we must first ascertain whether our system is observable at equilibrium. To show this, we linearize the observation model for the task of hovering. Note that the gyroscope provides a low-noise measurement of angular velocity $\bm{\omega}$. In cases like this in which the sensor provides more precise information than can be derived knowing the system’s actuated inputs (e.g. torque applied by the aircraft’s rotors or wings), it is common practice to reformulate the Kalman Filter to use the sensor’s readings as “input” $\bm{u}$, and to ignore actual system input. This has an additional advantage that one less state variable is required for each such measurement. Our system includes gyroscope readings as one of its inputs and hence angular velocity $\omega$ does not appear among the state variables. Our observation model for the suite of sensors is:
@@
where $\Omega_{m}$ is the optic flow measurements (rad/s) from optic flow sensor, $z_{m}$ is the measurement (m) from the pressure sensor, and $n_{o}$ and $n_{p}$ are the noise terms for the two sensors. We assume noise terms are zero-mean Gaussian white noise, a standard assumption in sensor noise modeling to simplify the analysis. The measurement model for the optic flow in the $x$ direction, as measured by the optic flow camera is a nonlinear function of the state [6]:
@@
where $\omega_{m}$ is the angular velocity measurement (rad/s) from gyroscope. The optic flow measurement depends on $z$. We can take the Jacobian linearization at a desired altitude $z_{d}$:
@@
With this linearization, our measurement model can be cast into the Kalman Filter model, $\bm{y}=C\bm{q}+D\bm{u}+\bm{n}$, where
@@
The Kalman Filter observer estimates $\hat{\bm{q}}$ by numerically integrating the system:
@@
We calculated optic flow variance $R_{n1}$ to be 0.017 by subtracting its value from that predicted by mocap using Eq. (3) during a constant-velocity horizontal translational flight. We calculated the variance of the pressure sensor $R_{n2}$ to be 0.0055 by taking measurements while holding it at a constant altitude for 60 seconds:
$R_{n}=\mathrm{diag}(R_{n1},R_{n2})=\mathrm{diag}(0.017,0.0055).$
To estimate the process noise covariance matrix $Q_{n}$, we employed a hybrid optimization approach, combining grid search for coarse parameter initialization, gradient descent for fine-tuning, and manual adjustment to minimize the RMSE between TinySense state estimates and mocap ground truth:
$Q_{n}=\mathrm{diag}(0.0124^{2},0.001^{2},0.22^{2}).$
For the process noise (disturbance) matrix $G$, we assumed that the noise enters the system as white noise disturbance adding to the derivatives of the three states $\theta$, $v_{x}$, and $z$. Based on these, the Kalman gain $K$ was computed using the lqe command in python-control [36]:
$G=\left[\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\end{matrix}\right],\quad K=\left[\begin{matrix}0.095&0\\ -1.32&0\\ 0&3\end{matrix}\right].$
@@
Figure 4: Sensor measurements from the onboard gyroscope, optic flow sensor, and pressure sensor of TinySense are compared with those from the Crazyflie for three different flight experiments (a), (b), and (c); trajectories are described in the main text and corresponding state estimates in Fig. 5. In the altitude plots, altitude is measured by the Crazyflie’s laser rangefinder, while the TinySense uses a pressure sensor; mocap data is also included for comparison.